Find the velocity at which the mass of a particle is doubleits rest mass

Grade:12th pass

1 Answers

Arun
5 years ago
Dear student
 
2 m c^2 = m c^2 / sqrt (1 – v^2/ c^2)
 
1- v^2 /c^2 = ¼
 
v = sqrt(3) * c /2 
 
hope it helps
 
Regards
Arun